(4x^4 z^2)^5 (x^3 yz^2)^4
Found 2 solutions by stanbon, mandymae:
Answer by stanbon(75887)

(27)^(x-2)=(9^1/3)^x
3^(3x-6) = 3^(2x/3)
The logs(exponents) must be equal:
3x-6 = (2/3)x
9x-18 = 2x
7x = 18
x = 18/7
--------------
You are correct.
Cheers,
Stan H.

Answer by mandymae(2)